x + 2(xp − y) + p^2 = 0


1
Expert's answer
2021-11-04T20:01:18-0400

Solution;

"x+2(xp-y)+p^2=0"

"x+2xp-2y+p^2=0"

"2y=x+2xp+p^2" .....(1)

"2y=x(1+2p)+p^2" ....(1a)

Differentiate (1) with respect forms;

"2\frac{dy}{dx}=2p=1+2p+2x\frac{dp}{dx}+2p\frac{dp}{dx}"

Resolve as;

"0=1+2\frac{dp}{dx}(x+p)"

"\frac{dp}{dx}=\frac{-1}{2(x+p)}"

"\frac{dx}{dp}=-2(x+p)"

"\frac{dx}{dp}+2x=-2p"

The I.F;

"I.F=e^{\int{2}dp}=e^{2p}"

Hence;

"x×I.F=\int-2pe^{2p}"

"xe^{2p}=\frac{-(2p-1)e^{2p}}{2}+c"

Therefore;

"x=\frac{1-2p}{2}+\frac{c}{e^{2p}}" ....(3)

Substitute (2) into (1a);

"2y=[\frac{1-2p}{2}+\frac{c}{e^{2p}}](1+2p)+p^2"

"y=[\frac{1-2p}{2}+\frac{c}{e^{2p}}]\frac{(1+2p)}{2}+\frac{p^2}{2}"